\begin{thebibliography}{10}

\bibitem{art1}
How {Many} {Photos} {Will} be {Taken} in 2020?

\bibitem{art2}
28+ {Selfie} {Statistics}, {Demographics}, \& {Fun} {Facts} (2022), March 2022.

\bibitem{Fang2014}
Yuming Fang, Kai Zeng, Zhou Wang, Weisi Lin, Zhijun Fang, and Chia-Wen Lin.
\newblock Objective quality assessment for image retargeting based on
  structural similarity.
\newblock {\em IEEE Journal on Emerging and Selected Topics in Circuits and
  Systems}, 4(1):95--105, 2014.

\bibitem{Gao2009}
Xinbo Gao, Wen Lu, Dacheng Tao, and Xuelong Li.
\newblock Image quality assessment based on multiscale geometric analysis.
\newblock {\em IEEE Transactions on Image Processing}, 18(7):1409--1423, 2009.

\bibitem{Golestaneh2014}
S~Alireza Golestaneh and Damon~M Chandler.
\newblock No-reference quality assessment of jpeg images via a quality
  relevance map.
\newblock {\em IEEE signal processing letters}, 21(2):155--158, 2014.

\bibitem{Gu2016}
Ke~Gu, Shiqi Wang, Guangtao Zhai, Siwei Ma, Xiaokang Yang, Weisi Lin, Wenjun
  Zhang, and Wen Gao.
\newblock Blind quality assessment of tone-mapped images via analysis of
  information, naturalness, and structure.
\newblock {\em IEEE Transactions on Multimedia}, 18(3):432--443, 2016.

\bibitem{Gu2014}
Ke~Gu, Guangtao Zhai, Xiaokang Yang, and Wenjun Zhang.
\newblock An efficient color image quality metric with local-tuned-global
  model.
\newblock In {\em Image Processing (ICIP), 2014 IEEE International Conference
  on}, pages 506--510. IEEE, 2014.

\bibitem{Gu2015}
Ke~Gu, Guangtao Zhai, Xiaokang Yang, and Wenjun Zhang.
\newblock Using free energy principle for blind image quality assessment.
\newblock {\em IEEE Transactions on Multimedia}, 17(1):50--63, 2015.

\bibitem{Kundu2015}
Debarati Kundu and Brian~L Evans.
\newblock Full-reference visual quality assessment for synthetic images: A
  subjective study.
\newblock In {\em Image Processing (ICIP), 2015 IEEE International Conference
  on}, pages 2374--2378. IEEE, 2015.

\bibitem{Larson2010}
Eric~C Larson and Damon~M Chandler.
\newblock Most apparent distortion: full-reference image quality assessment and
  the role of strategy.
\newblock {\em Journal of Electronic Imaging}, 19(1):011006--011006, 2010.

\bibitem{Li2016}
Leida Li, Wenhan Xia, Yuming Fang, Ke~Gu, Jinjian Wu, Weisi Lin, and Jiansheng
  Qian.
\newblock Color image quality assessment based on sparse representation and
  reconstruction residual.
\newblock {\em Journal of Visual Communication and Image Representation},
  38:550--560, 2016.

\bibitem{Li2015}
Leida Li, Yu~Zhou, Jinjian Wu, Weisi Lin, and Haoliang Li.
\newblock Gridsar: Grid strength and regularity for robust evaluation of
  blocking artifacts in jpeg images.
\newblock {\em Journal of Visual Communication and Image Representation},
  30:153--163, 2015.

\bibitem{Liang2010}
Luhong Liang, Shiqi Wang, Jianhua Chen, Siwei Ma, Debin Zhao, and Wen Gao.
\newblock No-reference perceptual image quality metric using gradient profiles
  for jpeg2000.
\newblock {\em Signal Processing: Image Communication}, 25(7):502--516, 2010.

\bibitem{Lin2011}
Weisi Lin and C-C~Jay Kuo.
\newblock Perceptual visual quality metrics: A survey.
\newblock {\em Journal of Visual Communication and Image Representation},
  22(4):297--312, 2011.

\bibitem{Liu2012}
Anmin Liu, Weisi Lin, and Manish Narwaria.
\newblock Image quality assessment based on gradient similarity.
\newblock {\em IEEE Transactions on Image Processing}, 21(4):1500--1512, 2012.

\bibitem{Mittal2012}
Anish Mittal, Anush~Krishna Moorthy, and Alan~Conrad Bovik.
\newblock No-reference image quality assessment in the spatial domain.
\newblock {\em IEEE Transactions on Image Processing}, 21(12):4695--4708, 2012.

\bibitem{Nafchi2016}
Hossein~Ziaei Nafchi, Atena Shahkolaei, Rachid Hedjam, and Mohamed Cheriet.
\newblock Mug: A parameterless no-reference jpeg quality evaluator robust to
  block size and misalignment.
\newblock {\em IEEE Signal Processing Letters}, 23(11):1577--1581, 2016.

\bibitem{Narwaria2010}
Manish Narwaria and Weisi Lin.
\newblock Objective image quality assessment based on support vector
  regression.
\newblock {\em IEEE Transactions on Neural Networks}, 21(3):515--519, 2010.

\bibitem{Narwaria2012}
Manish Narwaria, Weisi Lin, Ian~Vince McLoughlin, Sabu Emmanuel, and Liang-Tien
  Chia.
\newblock Fourier transform-based scalable image quality measure.
\newblock {\em IEEE Transactions on Image Processing}, 21(8):3364--3377, 2012.

\bibitem{Ponomarenko2013}
Nikolay Ponomarenko, Oleg Ieremeiev, Vladimir Lukin, Karen Egiazarian, Lina
  Jin, Jaakko Astola, Benoit Vozel, Kacem Chehdi, Marco Carli, Federica
  Battisti, et~al.
\newblock Color image database tid2013: Peculiarities and preliminary results.
\newblock In {\em Visual Information Processing (EUVIP), 2013 4th European
  Workshop on}, pages 106--111. IEEE, 2013.

\bibitem{Ponomarenko2009}
Nikolay Ponomarenko, Vladimir Lukin, Alexander Zelensky, Karen Egiazarian,
  Marco Carli, and Federica Battisti.
\newblock Tid2008-a database for evaluation of full-reference visual quality
  assessment metrics.
\newblock {\em Advances of Modern Radioelectronics}, 10(4):30--45, 2009.

\bibitem{Reza1961}
Fazlollah~M Reza.
\newblock {\em An introduction to information theory}.
\newblock Inc, New York, 1994.

\bibitem{Sheikh2006}
Hamid~R Sheikh, Muhammad~F Sabir, and Alan~C Bovik.
\newblock A statistical evaluation of recent full reference image quality
  assessment algorithms.
\newblock {\em IEEE Transactions on image processing}, 15(11):3440--3451, 2006.

\bibitem{Tan2000}
KT~Tan and Mohammed Ghanbari.
\newblock Blockiness detection for mpeg2-coded video.
\newblock {\em IEEE Signal Processing Letters}, 7(8):213--215, 2000.

\bibitem{Tan2000a}
KT~Tan and Mohammed Ghanbari.
\newblock Frequency domain measurement of blockiness in mpeg-2 coded video.
\newblock In {\em Image Processing, 2000. Proceedings. 2000 International
  Conference on}, volume~3, pages 977--980. IEEE, 2000.

\bibitem{Tao2009}
Dacheng Tao, Xuelong Li, Wen Lu, and Xinbo Gao.
\newblock Reduced-reference iqa in contourlet domain.
\newblock {\em IEEE Transactions on Systems, Man, and Cybernetics, Part B
  (Cybernetics)}, 39(6):1623--1627, 2009.

\bibitem{Wang2011a}
Shiqi Wang, Abdul Rehman, Zhou Wang, Siwei Ma, and Wen Gao.
\newblock Ssim-inspired divisive normalization for perceptual video coding.
\newblock In {\em Image Processing (ICIP), 2011 18th IEEE International
  Conference on}, pages 1657--1660. IEEE, 2011.

\bibitem{Wang2012}
Shiqi Wang, Abdul Rehman, Zhou Wang, Siwei Ma, and Wen Gao.
\newblock Ssim-motivated rate-distortion optimization for video coding.
\newblock {\em IEEE Transactions on Circuits and Systems for Video Technology},
  22(4):516--529, 2012.

\bibitem{Wang2013}
Shiqi Wang, Xiang Zhang, Siwei Ma, and Wen Gao.
\newblock Reduced reference image quality assessment using entropy of
  primitives.
\newblock In {\em Picture Coding Symposium (PCS), 2013}, pages 193--196. IEEE,
  2013.

\bibitem{Wang2002a}
Z~Wang and Alan~C Bovik.
\newblock Image and multidimensional signal processing-a universal image
  quality index.
\newblock {\em IEEE Signal Processing Letters}, 9(3):81--84, 2002.

\bibitem{Wang2000}
Zhou Wang, Alan~C Bovik, and BL~Evan.
\newblock Blind measurement of blocking artifacts in images.
\newblock In {\em Image Processing, 2000. Proceedings. 2000 International
  Conference on}, volume~3, pages 981--984. Ieee, 2000.

\bibitem{Wang2004}
Zhou Wang, Alan~C Bovik, Hamid~R Sheikh, and Eero~P Simoncelli.
\newblock Image quality assessment: from error visibility to structural
  similarity.
\newblock {\em IEEE transactions on image processing}, 13(4):600--612, 2004.

\bibitem{Wang2002}
Zhou Wang, Hamid~R Sheikh, and Alan~C Bovik.
\newblock No-reference perceptual quality assessment of jpeg compressed images.
\newblock In {\em Image Processing. 2002. Proceedings. 2002 International
  Conference on}, volume~1, pages I--I. IEEE, 2002.

\bibitem{Wu1997}
HR~Wu and M~Yuen.
\newblock A generalized block-edge impairment metric for video coding.
\newblock {\em IEEE Signal Processing Letters}, 4(11):317--320, 1997.

\bibitem{Wu2013}
Jinjian Wu, Weisi Lin, Guangming Shi, and Anmin Liu.
\newblock Perceptual quality metric with internal generative mechanism.
\newblock {\em IEEE Transactions on Image Processing}, 22(1):43--54, 2013.

\bibitem{Zhang2014}
Lin Zhang, Ying Shen, and Hongyu Li.
\newblock Vsi: A visual saliency-induced index for perceptual image quality
  assessment.
\newblock {\em IEEE Transactions on Image Processing}, 23(10):4270--4281, 2014.

\bibitem{Zhang2015}
Lin Zhang, Lei Zhang, and Alan~C Bovik.
\newblock A feature-enriched completely blind image quality evaluator.
\newblock {\em IEEE Transactions on Image Processing}, 24(8):2579--2591, 2015.

\bibitem{Zhang2011}
Lin Zhang, Lei Zhang, Xuanqin Mou, and David Zhang.
\newblock Fsim: A feature similarity index for image quality assessment.
\newblock {\em IEEE transactions on Image Processing}, 20(8):2378--2386, 2011.

\end{thebibliography}

\documentclass{article}
\usepackage{cite}
\usepackage{url}
\usepackage[T1]{fontenc}
\usepackage[latin9]{inputenc}
\usepackage{color}
\usepackage{array}
\usepackage{mathtools}
\usepackage{multirow}
\usepackage{amssymb}
\usepackage{setspace}
\usepackage{array}
\date{}
\newcolumntype{L}{>{\centering\arraybackslash}m{3cm}}

\makeatletter

\begin{document}

\title{No reference image quality assessment metric based on regional mutual
information among images}

\newcommand*\samethanks[1][\value{footnote}]{\footnotemark[#1]}
\author{Vinay Kumar\thanks{Department of Electronics \& Communication Engineering, Thapar Institute of Engineering and Technology,
Patiala}
\and Vivek Singh Bawa\thanks{Visual Artificial Intelligence Laboratory, Oxford Brookes University, UK.}
\and Rahul Upadhyay\samethanks[1]}

\maketitle
\begin{abstract}
With the inclusion of camera in daily life, an automatic no reference
image quality evaluation index is required for automatic classification
of images. The present manuscripts proposes a new No Reference Regional
Mutual Information based technique for evaluating the quality of an
image. We use regional mutual information on subsets of the complete
image. Proposed technique is tested on four benchmark natural image
databases, and one benchmark synthetic database. A comparative analysis
with classical and state-of-art methods indicate superiority of the
present technique for high quality images and comparable for other
images of the respective databases. 
\end{abstract}

\section{Introduction}

With the advent of inexpensive and good quality mobile cameras storage,
transmission and compression of images has become a standard practice
among technical and non technical masses. Large number of people have
mobile phones with camera capturing trillions of photographs every
year \cite{art1}, approximately 24 billion selfies were uploaded
to Google in year 2015 \cite{art2} and increasing exponentially with
every passing year. Unlimited space for uploading images on Google
photos (and large space on other web servers; for example, Flickr,
Pinterest, etc) facilitates and influences people to capture many
photographs of the same situation. Searching the good quality images
from this ever (exponentially) increasing large quantity is impossible
task for a human being. Therefore, it becomes pertinent to design
and develop better automatic and no-reference image quality assessment
system. These systems will help; for example, in evaluating the image
information and (possibly) retain the best out of plethora, find out
the quality in real time, selecting camera settings for best results,
etc. This drives researchers to develop better auto no reference image
quality measurement techniques \cite{Lin2011,Liu2012,Narwaria2010,Narwaria2012,Liang2010,Wang2013,Wang2002a}. 

Researchers generally talk about three types of image quality assessment
(IQA) techniques: 
\begin{enumerate}
\item full reference \cite{Zhang2011,Li2016,Wu2013,Gu2014,Wang2012,Wang2011a,Zhang2014,Fang2014}, 
\item reduced reference IQA \cite{Tao2009,Gao2009}, and 
\item no-reference IQA \cite{Mittal2012,Gu2015,Gu2016,Zhang2015}. 
\end{enumerate}
First type of IQA assumes that human beings are sensitive to degradations,
second indicates that we are more sensitive to few key features extracted
from the image. The current proposed technique lies in the last category. 

Various techniques for objective image quality measurement are discussed
in literature. Since human visual system is a complex set of decision
making processes, available IQA methods are still not as good as the
human visual decisions. We discuss some of the relevant and prevailing
methods in rest of the present section.

Wu et al \cite{Wu1997} used measurement of blocking effect in horizontal
and vertical directions and differences at block boundaries in horizontal
and vertical directions, respectively. Tan et al \cite{Tan2000a}
analyzes magnitude and phase information in a harmonics to measure
the quality of the image. Another \cite{Tan2000} model was developed
for measuring block effects in an image. Wang et al \cite{Wang2000}
used energy based measurements to find the blocking artifacts in an
image. These blocking effects become fundamental building blocks for
measurement of quality of an image.

While transmitting or storing, image quality (IQ) measurement plays
a crucial role to evaluate and choose the correct image. The ultimate
goal of IQ measurement is assigning a quantitative value to perception
to human observers. Researchers perform this task with the help of
crowd sourcing and acquiring Mean Opinion Score (MOS). MOS or its
modified versions compared with the IQA values become basis for quality
of the IQA index. 

In the next section we discuss proposed method followed, in section
\ref{sec:Experimental-Results}, by experimental results and discussion.
We close the manuscript with conclusion and references. 

\section{\label{sec:Methodology}Methodology}

The proposed index No Reference Regional Mutual Information (NrMI)
predicts the quality of an image with the help of following procedure. 

Given an image matrix $\Phi(x,y)\in\mathbb{Z}^{n\times m}$. Another
version of matrix $\Phi(x,y)$ is created and is depicted by $\Phi'_{\theta}(x',y')$,
where

\begin{equation}
\left[\begin{array}{c}
x'\\
y'
\end{array}\right]=\left[\begin{array}{cc}
cos\theta & -sin\theta\\
sin\theta & cos\theta
\end{array}\right]\left[\begin{array}{c}
x\\
y
\end{array}\right]\label{eq:thetaRot}
\end{equation}

To make $\Phi$ and $\Phi'_{\theta}$ of same size equations \ref{eq:vect}
and \ref{eq:invvect} are applied. 
\begin{equation}
\Phi_{vec}(v)=vec(\Phi'_{\theta}(x',y'))\label{eq:vect}
\end{equation}

\begin{equation}
\Phi_{\theta v}(x,y)=vec^{-1}(\Phi_{vec}(v))\vcentcolon\Phi_{\theta}(x,y)\in\mathbb{R}^{ab}\rightarrow\mathbb{R}^{n\times m}\label{eq:invvect}
\end{equation}

We divide $\Phi(x,y)$ into disjoint group of $n$ sub-matrices, $\eta_{k}^{a\times b}\colon k\in\mathbb{Z^{>}}$where
$\eta_{k}:\eta_{k}\subset\Phi$. $\eta_{k}$ contains $q$ member
of perfect subsets of $\Phi$ (such that $k/m\in\mathbb{Z^{>}}$),
for it is obvious that if a subset is perfect, then there is no information
loss. Every element of $\eta_{k}$ represents a segment of original
image $\Phi$. $\Phi'_{\theta}(x',y')$ is divided into sub-matrices
$\eta_{k,\theta}^{a\times b}$ ($\equiv\eta_{k}$). 

We choose size of $\eta_{k}$ (and consequently $\eta_{k,\theta}$)
to be $3\times3$, which makes sure that the values within $\eta_{k}$
will not be varying significantly except when sub-matrix lies at an
edge in $\Phi(x,y)$ (or $\Phi'_{\theta}(x',y')$). The value of $\theta$
is $\pi/2$, one can choose any value for $\theta$ but $\pi/2$ provides
maximum shift, and equation \ref{eq:thetaRot} is rewritten as

\begin{equation}
\left[\begin{array}{c}
x'\\
y'
\end{array}\right]=\left[\begin{array}{c}
-y\\
x
\end{array}\right]
\end{equation}

which relates image matrices $\Phi$ and $\Phi'_{\theta}$ by equation
\ref{eq:phiRel} 
\begin{equation}
\Phi(x,y)=\Phi'_{\theta}(-y,x)\label{eq:phiRel}
\end{equation}

Let sub-matrices $\eta_{k}$ (and $\eta_{k,\theta}$) be represented
by 

\begin{equation}
\begin{array}{ccc}
\eta_{k} & = & \left[\begin{array}{ccc}
e_{11} & e_{12} & e_{13}\\
e_{21} & e_{22} & e_{23}\\
e_{31} & e_{32} & e_{33}
\end{array}\right]\\
\eta_{k,\theta} & = & \left[\begin{array}{ccc}
e_{11,\theta} & e_{12,\theta} & e_{13,\theta}\\
e_{21,\theta} & e_{22,\theta} & e_{23,\theta}\\
e_{31,\theta} & e_{32,\theta} & e_{33,\theta}
\end{array}\right]
\end{array}\label{eq:elements}
\end{equation}

From matrices of equation \ref{eq:elements} calculate matrix $M_{e}$ 

\begin{equation}
M_{e}=\left[e_{11}\;e_{12}\;e_{13}\;e_{11,\theta}\;e_{12,\theta}\;e_{13,\theta}\;e_{21}\ldots e_{23}\;e_{21,\theta}\;\ldots\;e_{33,\theta}\right]\label{eq:MatrixP}
\end{equation}

Center the values at the origin and represent it by $M_{e,0}$ by
\begin{equation}
M_{e,0}=M_{e}-\frac{1}{N}\sum_{i}^{N}p_{i}
\end{equation}

where $p_{i}$ are elements of matrix $M_{e}$ and $N=9+9=18$.

Find covariance 
\begin{equation}
C=\frac{1}{N}M_{e,0}M_{e,0}^{T}
\end{equation}

Estimate joint entropy \footnote{\label{fn:Joint-and-marginal}Joint and marginal entropy is given
by \cite{Reza1961}
\begin{equation}
H_{g}(\Sigma_{d})=log((2\pi e)^{\frac{d}{2}}det(\Sigma_{d})^{\frac{1}{2}})
\end{equation}
which represents the entropy of a normally distributed set of points
in $\Re^{d}$ with covariance matrix $\Sigma_{d}$.}
\[
H_{g}(C)
\]

Estimate marginal entropy\ref{fn:Joint-and-marginal} $H_{g}(C_{A})$
and $H_{g}(C_{B})$, where $C_{A}$ and $C_{B}$ are top left and
bottom right $\frac{d}{2}\times\frac{d}{2}$ matrices of $C$.  $d$
is a relationship defined as

\begin{equation}
d=2(2r+1)^{2}\label{eq:val-r}
\end{equation}

where, $r$ is the size of sub matrix under consideration; that is,
the size of MB for which we are going to calculate the similarity
within the matrix; for example  in Figure value of $r$ is $1$. 

Calculate Regional Mutual Information
\begin{equation}
M_{rmi}=H_{g}(C_{A})+H_{g}(C_{B})-H_{g}(C)\label{eq:RMI}
\end{equation}

$M_{rmi}$ gives a measure of regional mutual information between
$\Phi(x,y)$ and $\Phi_{\theta}(x',y')$. 

A weight function for RMI is calculated with equation of $\Phi_{vec}(v)$
is calculated next 
\begin{equation}
\Phi_{wg}=\mathbf{E}\left[(\Phi_{vec}(v)-\mathbf{E}[\Phi_{vec}])^{2}\right]\label{eq:varweight}
\end{equation}

The relative quality of an image is given by
\begin{equation}
NrMI_{i}=M_{rmi,i}*\Phi_{wg,i}\label{eq:qualityMat}
\end{equation}

where $i\in i^{th}$ image in the image sequence.

\section{\label{sec:Experimental-Results}Experimental Results }

In this section we validate our method through application on various
benchmark state-of-art and classical databases. Experiments are conducted
with five standard databases of natural and one of synthetic images.
The natural image databases are TID 2008 \cite{Ponomarenko2009} with
1699 images, TID 2013 \cite{Ponomarenko2013} with 2483 images, CID
2013 \cite{Larson2010}, LIVE \cite{Sheikh2006}, MEFD with 550 images
each. While ESPL \cite{Kundu2015}, a database consisting of 550 synthetic
images, is used for evaluation of the current algorithm.

For objective evaluation SRCC (Spearman's Rank Correlation Coefficient)
and PLCC (Pearson Linear Correlation Coefficient) matrices are used.
These metrics give a measure of prediction monotonicity and linearity,
respectively. 

Table \ref{tab:Performance-Comparison-of} presents a comparative
view of various index of quantitative quality measures. Blue color
values in table \ref{tab:Performance-Comparison-of} indicate best
results. Since no-reference quality measurement system requires complex
set of interdependent parameters to work as efficiently as human beings,
therefore every system has certain advantage over others under certain
conditions. From the table it becomes clear that proposed method evaluates
the images better than SSIM for most of the databases. Since the proposed
method uses underlying regional geometric information by splitting
the set into disjoint group of sub-sets; therefore every small change
in geometry (including presence of undetectable noise for human visual
system) changes the qualitative measure. 

Specifically with images of high quality (databases MEFD and ESPL)
the proposed method performs much better than SSIM \cite{Wang2004}
and other state-of-art techniques. Since proposed technique considers
underlying geometry of the image, high quality images distorted by
small amount of noise produce lower value of the quality index. This
lower value in turn will be helpful to take corrective measures to
develop noise removal or better compression algorithms. 

\begin{table*}
\centering

\begin{tabular}{|>{\centering}p{2cm}|>{\centering}p{2cm}|>{\centering}p{1cm}|>{\centering}p{1cm}|>{\centering}p{1cm}|>{\centering}p{1cm}|>{\centering}p{1cm}|>{\centering}p{1cm}|>{\centering}p{1cm}|}
\hline 
{\scriptsize{}Database} & {\scriptsize{}Statistical Measurement} & {\scriptsize{}SSIM }\cite{Wang2004}{\scriptsize{} } & \multirow{1}{1cm}{{\scriptsize{}NR \cite{Wang2002}}} & {\scriptsize{}NJQA \cite{Golestaneh2014}} & {\scriptsize{}NR \cite{Li2015}} & {\scriptsize{}MUG NR \cite{Nafchi2016}} & {\scriptsize{}MUG$^{+}$ NR \cite{Nafchi2016}} & {\scriptsize{}PM}\tabularnewline
\hline 
\hline 
{\scriptsize{}TID 2008 \cite{Ponomarenko2009}} & {\scriptsize{}PLCC} & \textcolor{blue}{\scriptsize{}0.954} & {\scriptsize{}0.952} & {\scriptsize{}0.944} & \textcolor{black}{\scriptsize{}0.951} & {\scriptsize{}0.941} & \textcolor{black}{\scriptsize{}0.953} & {\scriptsize{}0.868}\tabularnewline
\hline 
 & {\scriptsize{}SRCC} & \textcolor{blue}{\scriptsize{}0.925} & {\scriptsize{}0.913} & {\scriptsize{}0.8993} & {\scriptsize{}0.917} & \textcolor{black}{\scriptsize{}0.917} & \textcolor{black}{\scriptsize{}0.924} & {\scriptsize{}0.832}\tabularnewline
\hline 
{\scriptsize{}TID 2013 \cite{Ponomarenko2013}} & {\scriptsize{}PLCC} & \textcolor{black}{\scriptsize{}0.954} & {\scriptsize{}0.953} & {\scriptsize{}0.948} & \textcolor{black}{\scriptsize{}0.955} & {\scriptsize{}0.942} & \textcolor{blue}{\scriptsize{}0.955} & {\scriptsize{}0.887}\tabularnewline
\hline 
 & {\scriptsize{}SRCC} & {\scriptsize{}0.9200} & \textcolor{black}{\scriptsize{}0.927} & {\scriptsize{}0.886} & \textcolor{blue}{\scriptsize{}0.931} & {\scriptsize{}0.908} & {\scriptsize{}0.919} & {\scriptsize{}0.842}\tabularnewline
\hline 
{\scriptsize{}CID 2013 \cite{Larson2010}} & {\scriptsize{}PLCC} & \textcolor{black}{\scriptsize{}0.979} & \textcolor{black}{\scriptsize{}0.975} & {\scriptsize{}0.954} & \textcolor{blue}{\scriptsize{}0.979} & {\scriptsize{}0.9679} & {\scriptsize{}0.972} & {\scriptsize{}0.789}\tabularnewline
\hline 
 & {\scriptsize{}SRCC} & {\scriptsize{}0.955} & \textcolor{black}{\scriptsize{}0.955} & {\scriptsize{}0.925} & \textcolor{blue}{\scriptsize{}0.957} & {\scriptsize{}0.930} & {\scriptsize{}0.937} & {\scriptsize{}0.798}\tabularnewline
\hline 
{\scriptsize{}LIVE \cite{Sheikh2006}} & {\scriptsize{}PLCC} & \textcolor{blue}{\scriptsize{}0.979} & \textcolor{black}{\scriptsize{}0.979} & {\scriptsize{}0.956} & {\scriptsize{}0.976} & {\scriptsize{}0.965} & {\scriptsize{}0.973} & {\scriptsize{}0.962}\tabularnewline
\hline 
 & {\scriptsize{}SRCC} & {\scriptsize{}0.946} & \textcolor{blue}{\scriptsize{}0.974} & {\scriptsize{}0.956} & \textcolor{black}{\scriptsize{}0.973} & {\scriptsize{}0.959} & {\scriptsize{}0.968} & {\scriptsize{}0.959}\tabularnewline
\hline 
{\scriptsize{}ESPL \cite{Kundu2015}} & {\scriptsize{}PLCC} & {\scriptsize{}0.943} & {\scriptsize{}0.960} & {\scriptsize{}0.809} & \textcolor{blue}{\scriptsize{}0.962} & {\scriptsize{}0.940} & {\scriptsize{}0.937} & \textcolor{blue}{\scriptsize{}0.962}{\scriptsize{} }\tabularnewline
\hline 
 & {\scriptsize{}SRCC} & {\scriptsize{}0.904} & {\scriptsize{}0.933} & {\scriptsize{}0.739} & \textcolor{black}{\scriptsize{}0.933} & {\scriptsize{}0.928} & {\scriptsize{}0.927} & \textcolor{blue}{\scriptsize{}0.959}\tabularnewline
\hline 
\end{tabular}

\caption{\label{tab:Performance-Comparison-of}Performance comparison of no
reference image quality measure for TID 2008 \cite{Ponomarenko2009},
TID 2013 \cite{Ponomarenko2013}, CID 2013 \cite{Larson2010}, LIVE
\cite{Sheikh2006}, MEFD, and ESPL \cite{Kundu2015} databases. Blue
color values represent best performing technique in terms of corresponding
SRCC and PLCC statistical measures. }
\end{table*}

\section{\label{sec:Conclusion}Conclusion}

Present manuscript investigates the problem of no-reference quality
assessment. A novel technique has been proposed for the assessment
based on underlying geometry of the image. The technique is applied
on various databases with different types of images. Results show
interesting trend and promising performance when compared with existing
literature. Since method utilizes mutual information approach it was
able to render better results for high quality images. 

In future we aim to study the effects of current technique by calculating
RMI on weighted image segments. The weights will be calculated based
on the importance of the region in the images, which in turn depends
on point of focus in human visual system. 

\bibliographystyle{plain}
\bibliography{myreferences}

\end{document}